OpenStudy (anonymous):

To convert 5 yards to inches, you would use the ratio 1yard/36 inches? True or False

hero (hero):

Would be the other way around

hero (hero):

36 in/1 yd

OpenStudy (anonymous):

tgyvbtgfv

OpenStudy (anonymous):

Its written 1yd/36in

OpenStudy (johnweldon1993):

False...if you do that...you'll get \[5 \space\text{yards} \times \frac{ 1 \space\text{yard} }{ 36\space\text{inches} }\] and that will become \[\frac{ 5\space\text{yards}^2 }{ 36\space\text{inches} }\] That is NOT what you want lol

OpenStudy (anonymous):

Lol ooh Tru

OpenStudy (johnweldon1993):

you would want it to be the other way around \[5\space\text{yards} \times \frac{ 36\space\text{inches} }{ 1\space\text{yard} }\]

OpenStudy (anonymous):

Alright so False right?

OpenStudy (johnweldon1993):

False is correct

OpenStudy (anonymous):

Thanxx

OpenStudy (johnweldon1993):

No problem!